# Mapping Cone Connections and their Yang-Mills Functional

### v1

@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
for $\gamma\in\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$. To denote the equivalence class, we use the bracket to denote a point $u=[\delta,g]\in P_{\rho}$. Note that $G$ acts on $P_{\rho}$ by $[\delta,g]h=[\delta,gh]$ for $h\in G$.
To define a connection on a principal bundle, recall there are two equivalent descriptions. We can define a connection on $P_{\rho}$ either as a horizontal subspace of $H_{u}\subset TP_{u}$ or as a one-form $A_{\rho}\in\Omega^{1}(P_{\rho},\mathfrak{g})$. They are related by $H_{u}=\ker A_{\rho}$. On $P_{\rho}$, we define a horizontal distribution by defining the horizontal lifts for any path $\alpha$ on $M$ as follows. Given an arbitrary point $[\delta,g]\in(P_{\rho})_{\alpha(0)}$ on the fiber of base point $\alpha(0)=\delta(1)\in M$, we define the horizontal lift of $\alpha$ starting from $[\delta,g]\in P_{\rho}$ by $\tilde{\alpha}(t)=[\alpha_{(t)}\delta,g]$ where $\alpha_{(t)}$ denotes the one-parameter family of paths within $\alpha$ starting at $\alpha(0)$ and ending at $\alpha(t)$. It is straightforward to check that this construction is well-defined and satisfies the $G$ invariance condition for a connection on $P_{\rho}$. We wil not need to explicitly write down $A_{\rho}$ in order to check that $A_{\rho}$ satisfies the cone-flat condition. We will express the curvature in terms of the holonomy group.
Step 2. Verify that the connection $(A_{\rho},B_{\rho})$ is cone-flat.
@@
The connection $(A_{\rho},B_{\rho})$ constructed above is cone-flat.
@@
Let $p\in M$ be an arbitrary point and $v_{1},v_{2}\in T_{p}M$ be arbitrary linearly independent vectors. Suppose $\zeta(v_{1},v_{2})=c\neq 0$. By Darboux’s theorem, we can find a local coordinate system $\{x_{1},\cdots,x_{m}\}$ such that $v_{1}=\frac{\partial}{\partial x_{1}}$, $v_{2}=\frac{\partial}{\partial x_{2}}$, and $\zeta=c\,dx_{1}\wedge dx_{2}+\bar{\zeta}$, where $\bar{\zeta}=dx_{3}\wedge dx_{4}+\ldots$ if $\dim M\geq 4$. Let $D_{t}$ be an infinitesimal parallelogram spanned by $\sqrt{t}v_{1}$ and $\sqrt{t}v_{2}$ in the local coordinate system, and $\gamma_{t}=\partial D_{t}$ be its boundary. Then we have $\int_{D_{t}}\zeta=c\,t$.
@@
Note that $\delta^{-1}\gamma_{t}\delta$ is contractible and its base point is $a$, i.e. $\delta^{-1}\gamma_{t}\delta\in\Psi^{a}_{a,0}(M)$. We can express $\delta^{-1}\gamma_{t}\delta=\exp(c\,t\,\xi)$, where $\xi$ is an element of the Lie algebra of $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ that generates the subgroup $\Psi^{a}_{a,0}(M)/\Psi^{a}_{a,\zeta}(M)\cong\mathbb{R}/H$ and $\int_{D_{t}}\zeta=\zeta(v_{1},v_{2})\,t=c\,t\in\mathbb{R}/H$. Note that $\xi$ is independent of the choice of $v_{1}$ and $v_{2}$. We thus find at $a\in M$
@@
where $d\rho$ maps the Lie algebra of $\Gamma$ into the $Lie(g)=\mathfrak{g}$. It is clear that $B_{\rho}$ as obtained above is covariantly constant as it is a constant when evaluated along horizontally lifted curves.
In the case where $\zeta(v_{1},v_{2})=0$ and $\dim M\geq 4$, then we can find a local coordinate system $\{x_{1},\cdots,x_{m}\}$ such that $v_{1}=\frac{\partial}{\partial x_{1}}$, $v_{2}=\frac{\partial}{\partial x_{3}}$, and $\zeta=dx_{1}\wedge dx_{2}+dx_{3}\wedge dx_{4}+\bar{\zeta}$, where $\bar{\zeta}$ is generated by other $dx_{i}\wedge dx_{j}$ locally. The proof goes through as above but with $c$ set to zero. ∎
@@
Suppose $\rho$ and $\bar{\rho}$ are conjugate homomorphisms from $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ to $G$. $(P_{\rho},A_{\rho})$ and $(P_{\bar{\rho}},A_{\bar{\rho}})$ are principal bundles with cone-flat connections constructed as above. Then $(P_{\rho},A_{\rho})$ and $(P_{\bar{\rho}},A_{\bar{\rho}})$ are equivalent.
@@
Suppose $\bar{\rho}=g_{0}\rho g_{0}^{-1}$, we define an automorphism on $\Psi_{a}(M)\times G$ by $(\delta,g)\mapsto(\delta,g_{0}g)$ and this automorphism induces an isomorphism $f:\Psi_{a}(M)\times_{\rho}G\to\Psi_{a}(M)\times_{\bar{\rho}}G$. For arbitrary path $\alpha$ on $M$ and $[\delta,g]_{\rho}\in\Psi_{a}(M)$ such that $\delta(1)=\alpha(0)$, the horizontal lift of $\alpha$ at $[\delta,g]_{\rho}$ is defined as $\tilde{\alpha}_{\rho}(t)=[\alpha_{(t)}\delta,g]_{\rho}$. On the other hand, the horizontal lift of $\alpha$ at $f\big{(}[\delta,g]_{\rho}\big{)}=[\delta,g_{0}g]_{\bar{\rho}}$ is defined as $\tilde{\alpha}_{\bar{\rho}}(t)=[\alpha_{(t)}\delta,g_{0}g]_{\bar{\rho}}$. So we have $\tilde{\alpha}_{\bar{\rho}}(t)=f\circ\tilde{\alpha}_{\rho}(t)$, which implies $f_{*}$ sends horizontal vectors to horizontal vectors. Thus, $f^{*}A_{\bar{\rho}}=A_{\rho}$. ∎
By this lemma, given any conjugacy classes $[\rho]$ of the morphisms from $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ to $G$, there is a corresponding $G$-bundle $P_{\rho}$ with a cone-flat connection $A_{\rho}$. It remains to show that this correspondence is bijective.
@@
Let $\rho,\bar{\rho}:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G$. $\rho$ and $\bar{\rho}$ are conjugate when $(P_{\rho},A_{\rho})$ and $(P_{\bar{\rho}},A_{\bar{\rho}})$ are equivalent.
@@
Suppose $f:P_{\rho}\to P_{\bar{\rho}}$ is a $G$-bundle isomorphism and $f^{*}A_{\bar{\rho}}=A_{\rho}$. Set $h:\Psi_{a}(M)\to G$ such that
@@
where $\delta\in\Psi_{a}(M)$ and $e$ is the identity of $G$. For each $\gamma\in\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$, we have
@@
For arbitrary $\delta\in\Psi_{a}(M)$, let $\delta_{(t)}$ denote the $t$-parametrized subpaths of $\delta$ starting at $\delta(0)=a$ and ending at $\delta(t)$. By the construction of the connections, $[\delta_{(t)},e]_{\rho}$ and $[\delta_{(t)},e]_{\bar{\rho}}$ are horizontal paths on $P_{\rho}$ and $P_{\bar{\rho}}\,$, respectively. The projection of these two paths on $M$ are exactly $\delta$. By the definition of $h$, we have
$f\big{(}[\delta_{(t)},e]_{\rho}\big{)}=[\delta_{(t)},e]_{\bar{\rho}}h(\delta_{(t)}).$
On the other hand, since $f^{*}A_{\bar{\rho}}=A_{\rho}$, $f\big{(}[\delta_{(t)},e]_{\rho}\big{)}$ is a horizontal lift of $\delta$ on $P_{\bar{\rho}}$ passing through $f\big{(}[\delta_{(0)},e]_{\rho}\big{)}$. As horizontal curves are equivariant by right $G$-action, $[\delta_{(t)},e]_{\bar{\rho}}h(\delta_{(0)})$ is also a horizontal lift of $\delta$ on $P_{\bar{\rho}}$ passing through $[\delta_{(0)},e]_{\bar{\rho}}h(\delta_{(0)})=f\big{(}[\delta_{(0)},e]_{\rho}\big{)}$. Therefore, we have
$f\big{(}[\delta_{(t)},e]_{\rho}\big{)}=[\delta_{(t)},e]_{\bar{\rho}}h(\delta_{(0)}).$
Comparing the two equations, we have $h(\delta_{(t)})=h(\delta_{(0)})$ for all $t\in[0,1]$. This implies $h(\delta)$ is equal to $h$ acting on the constant path at $a$. So $h$ is a constant. Therefore, we have $\rho$ and $\bar{\rho}$ are conjugate. ∎
@@
Suppose $\pi:P\to M$ be a principal $G$-bundle, and $A$ is a cone-flat connection on $P$ with curvature $F=\Phi\zeta$. There exists $\xi\in\mathfrak{g}$ such that for any contractible loop $\gamma$ starting at $a$ and any oriented disk $D$ in $M$ with $\partial D=\gamma$ ($\partial D$ and $\gamma$ also have the same orientation), the holonomy along $\gamma$ is
$hol(\gamma)=-\mathrm{exp}\Big{(}\int_{D}\zeta\cdot\xi\Big{)}.$
@@
Take a point $u_{0}\in P$ on the fiber of $a$. Consider the holonomy bundle
$\hat{P}=\{u\in P|\text{there exists horizontal path }\delta\text{ such that }\delta(0)=u_{0},\delta(1)=u\}.$
$\hat{P}$ is a principal bundle over $M$, and its structure group $\hat{G}$ is the holonomy group of $P$ at $u_{0}$. Let $\hat{A}$ be the restriction of $A$ on $\hat{P}$. By the holonomy theorem of Ambrose and Singer [AS], the Lie algebra of $\hat{G}$ is
$\hat{\mathfrak{g}}=\mathrm{span}\left\{F\left(v_{1}^{H},v_{2}^{H}\right)\big{|}v_{1}^{H},v_{2}^{H}\text{ are horizontal vectors at }u\text{ for some }u\in\hat{P}\right\}.$
By assumption, $F=\Phi\zeta$, or more precisely $F=\Phi(\pi^{*}\zeta)$. Since $\Phi$ is covariantly constant, it is equal to some $\xi\in\hat{\mathfrak{g}}$ at any point in $\hat{P}$. Hence, $F\left(v_{1}^{H},v_{2}^{H}\right)\in\mathbb{R}\xi$. So $\hat{\mathfrak{g}}$ is 1-dimensional and abelian. Then $\hat{A}=\xi\otimes\theta$ for some $\theta\in\Omega^{1}(M,Ad\,\hat{P})$ and $d\theta=\hat{\pi}^{*}\zeta$, where $\hat{\pi}:\hat{P}\to M$ is the projection.
For an arbitrary contractible loop $\gamma$ and a disk $D$ such that $\partial D=\gamma$, there exists a contractible neighborhood $U\subset M$ of $D$. Then $\hat{P}|_{U}=U\times\hat{G}$ is trivial. Let $\sigma:U\to\hat{P}|_{U}$ be a local section and $\psi:\hat{P}|_{U}\to U\times\hat{G}$ be a trivialization such that $\psi\circ\sigma(p)=(p,e)$ for $p\in M$. Then $(\psi^{-1})^{*}\hat{A}=(\xi\otimes\sigma^{*}\theta,0)+(0,MC_{\hat{G}})$, where $MC_{\hat{G}}:T\hat{G}\to\hat{\mathfrak{g}}$ is the Maurer-Cartan form of $\hat{G}$ sending a vector to the corresponding invariant vector field. Observe that $d(\sigma^{*}\theta)=\zeta$. The horizontal lift $\tilde{\gamma}$ of $\gamma$ with $\tilde{\gamma}(0)=\sigma(a)$ satisfies $\psi\circ\tilde{\gamma}(t)=(\gamma(t),g(t))$ with $g(t)\in\hat{G}$ and $g(0)=e$. Then
@@
$=(\sigma^{*}\theta)\big{(}\gamma^{\prime}(t)\big{)}\cdot\xi\Big{|}_{\big{(}\gamma(t_{0})\big{)}}+MC_{\hat{G}}\big{(}g^{\prime}(t)\big{)}\Big{|}_{\big{(}g(t_{0})\big{)}}.$
So we have
@@
$hol(\gamma)=g(1)=-\mathrm{exp}\Big{(}\int_{\gamma}\sigma^{*}\theta\cdot\xi\Big{)}=-\mathrm{exp}\Big{(}\int_{D}\zeta\cdot\xi\Big{)}.$
Now we prove the surjectivity.
@@
Let $\pi:P\to M$ be a principal $G$-bundle with a cone-flat connection $A$. There exists a morphism $\rho:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G$ such that $(P_{\rho},A_{\rho})$ and $(P,A)$ are equivalent.

### v2

@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
where $\gamma\in\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$. To denote the equivalence class, we use the bracket to denote a point $u=[\delta,g]\in P_{\rho}$. Note that $G$ acts on $P_{\rho}$ by $[\delta,g]h=[\delta,gh]$ for $h\in G$.
To define a connection on a principal bundle, recall that there are two ways to do so. We can define a connection on $P_{\rho}$ either as a horizontal subspace, $H_{u}\subset(TP_{\rho})_{u}$ at all $u\in P_{\rho}$, or as a one-form $A_{\rho}\in\Omega^{1}(P_{\rho},\mathfrak{g})$. They are related by $H_{u}=\ker A_{\rho}\big{|}_{u}$. On $P_{\rho}$, we define a horizontal distribution by defining the horizontal lifts for any path $\alpha$ on $M$ as follows. Given an arbitrary point $[\delta,g]\in(P_{\rho})_{\alpha(0)}$ on the fiber of base point $\alpha(0)=\delta(1)\in M$, we define the horizontal lift of $\alpha$ starting from $[\delta,g]\in P_{\rho}$ by $\tilde{\alpha}(t)=[\alpha_{(t)}\delta,g]$ where $\alpha_{(t)}$ denotes the one-parameter family of paths within $\alpha$ starting at $\alpha(0)$ and ending at $\alpha(t)$. It is straightforward to check that this construction is well-defined and satisfies the $G$ invariance condition for a connection on $P_{\rho}$. We will not need to explicitly write down $A_{\rho}$ as a one-form in order to check that $A_{\rho}$ satisfies the cone-flat condition. We will instead express the curvature in terms of the holonomy group.
Step 2. Verify that the connection $A_{\rho}$ is cone-flat, that is, $F_{A_{\rho}}=-\zeta B_{\rho}$ such that $B_{\rho}$ is covariantly constant.
@@
The connection $A_{\rho}$ constructed above is cone-flat.
@@
Let $p\in M$ be an arbitrary point and $v_{1},v_{2}\in T_{p}M$ be arbitrary linearly independent vectors. Suppose $\zeta(v_{1},v_{2})=c\neq 0$. By Darboux’s theorem, we can find a local coordinate system $\{x_{1},\ldots,x_{m}\}$ such that $v_{1}=\frac{\partial}{\partial x_{1}}$, $v_{2}=\frac{\partial}{\partial x_{2}}$, and $\zeta=c\,dx_{1}\wedge dx_{2}+\bar{\zeta}$, where $\bar{\zeta}=dx_{3}\wedge dx_{4}+\ldots,$ if $\dim M\geq 4$. Let $D_{t}$ be an infinitesimal parallelogram spanned by $\sqrt{t}v_{1}$ and $\sqrt{t}v_{2}$ in the local coordinate system, and $\gamma_{t}=\partial D_{t}$ be its boundary. Then we have $\int_{D_{t}}\zeta=c\,t$.
@@
Note that $\delta^{-1}\gamma_{t}\delta$ is contractible and its base point is $a\in M$, i.e. $\delta^{-1}\gamma_{t}\delta\in\Psi^{a}_{a,0}(M)$. We can express $\delta^{-1}\gamma_{t}\delta=\exp(c\,t\,\xi)$, where $\xi$ is an element of the Lie algebra of $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ that generates the subgroup $\Psi^{a}_{a,0}(M)/\Psi^{a}_{a,\zeta}(M)\cong\mathbb{R}/H$ and $\int_{D_{t}}\zeta=\zeta(v_{1},v_{2})\,t=c\,t\in\mathbb{R}/H$. Note that $\xi$ is independent of the choice of $v_{1}$ and $v_{2}$. We thus find at $p\in M$
@@
where $d\rho$ maps the Lie algebra of $\Gamma$ into $\mathfrak{g}=\operatorname{Lie}(G)$. It is clear that $B_{\rho}$ as obtained above is covariantly constant as it is a constant when evaluated along horizontally lifted curves.
Thus far, we have assumed $\zeta(v_{1},v_{2})\neq 0$. If however $\zeta(v_{1},v_{2})=0$ which may occur in $\dim M\geq 4$, then we can find a local coordinate system $\{x_{1},\ldots,x_{m}\}$ such that $v_{1}=\frac{\partial}{\partial x_{1}}$, $v_{2}=\frac{\partial}{\partial x_{3}}$, and $\zeta=dx_{1}\wedge dx_{2}+dx_{3}\wedge dx_{4}+\bar{\zeta}$, where $\bar{\zeta}$, if non-zero, is generated by other $dx_{i}\wedge dx_{j}$ locally. The proof goes through as above but with $c$ set to zero. ∎
@@
Suppose $\rho$ and $\bar{\rho}$ are conjugate homomorphisms from $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ to $G$, and also, $(P_{\rho},A_{\rho})$ and $(P_{\bar{\rho}},A_{\bar{\rho}})$ are principal bundles with cone-flat connections constructed as above. Then $(P_{\rho},A_{\rho})$ and $(P_{\bar{\rho}},A_{\bar{\rho}})$ are equivalent.
@@
Suppose $\rho$ and $\bar{\rho}$ are conjugate, i.e. $\bar{\rho}=g_{0}\rho g_{0}^{-1}$ for some $g_{0}\in G$. We consider the automorphism on $\Psi_{a}(M)\times G$ given by $(\delta,g)\mapsto(\delta,g_{0}g)$. This automorphism induces the desired bundle isomorphism
$f:\quad\Psi_{a}(M)\times_{\rho}G\quad\longrightarrow$ $\quad\quad\Psi_{a}(M)\times_{\bar{\rho}}G$
$(\delta\gamma,g)\sim(\delta,\rho(\gamma)g)\mapsto$ $\ \ (\delta\gamma,g_{0}g)\sim(\delta,\bar{\rho}(\gamma)g_{0}g)=(\delta,g_{0}\rho(\gamma)g)$
where the second line shows the map on the equivalence class for all $\gamma\in\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$.
Now to show $f^{*}A_{\bar{\rho}}=A_{\rho}$, let $[\delta,g]_{\rho}\in\Psi_{a}(M)\times_{\rho}G$ and $\alpha$ on $M$ be an arbitrary path starting at $\alpha(0)=\delta(1)$. As described in Step 1 in defining $A_{\rho}$, the horizontal lift of $\alpha$ starting at $[\delta,g]_{\rho}$ is defined as $\tilde{\alpha}_{\rho}(t)=[\alpha_{(t)}\delta,g]_{\rho}$. Likewise, the horizontal lift of $\alpha$ starting at $f\big{(}[\delta,g]_{\rho}\big{)}=[\delta,g_{0}g]_{\bar{\rho}}$ is defined as $\tilde{\alpha}_{\bar{\rho}}(t)=[\alpha_{(t)}\delta,g_{0}g]_{\bar{\rho}}$. Clearly, we have $\tilde{\alpha}_{\bar{\rho}}(t)=f\circ\tilde{\alpha}_{\rho}(t)$, which implies $f_{*}$ sends horizontal vectors to horizontal vectors as desired. ∎
By this lemma, given any conjugacy classes $[\rho]$ of the homomorphisms from $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ to $G$, there is a corresponding $G$-bundle $P_{\rho}$ with a cone-flat connection $A_{\rho}$. It remains to show that this correspondence is bijective.
@@
Let $\rho,\bar{\rho}:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G$. If $(P_{\rho},A_{\rho})$ and $(P_{\bar{\rho}},A_{\bar{\rho}})$ are equivalent, then $\rho$ and $\bar{\rho}$ are conjugate.
@@
Suppose $f:P_{\rho}\to P_{\bar{\rho}}$ is a $G$-bundle isomorphism and $f^{*}A_{\bar{\rho}}=A_{\rho}$. Let us define $h:\Psi_{a}(M)\to G$ such that
@@
where $\delta\in\Psi_{a}(M)$ and $e$ is the identity of $G$. Then, for each $\gamma\in\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$, we have
@@
For arbitrary $\delta\in\Psi_{a}(M)$, let $\delta_{(t)}$ denote the $t$-parametrized subpaths of $\delta$ starting at $\delta(0)=a$ and ending at $\delta(t)$. By the construction of the connections, $[\delta_{(t)},e]_{\rho}$ and $[\delta_{(t)},e]_{\bar{\rho}}$ are horizontal paths on $P_{\rho}$ and $P_{\bar{\rho}}\,$, respectively. The projection of these two paths on $M$ are exactly $\delta_{(t)}$. By the definition of $h$, we have
$f\big{(}[\delta_{(t)},e]_{\rho}\big{)}=[\delta_{(t)},e]_{\bar{\rho}}h(\delta_{(t)})\,.$ (4.4)
On the other hand, since $f^{*}A_{\bar{\rho}}=A_{\rho}$, $f\big{(}[\delta_{(t)},e]_{\rho}\big{)}$ is a horizontal lift of $\delta_{(t)}$ on $P_{\bar{\rho}}$ passing through $f\big{(}[\delta_{(0)},e]_{\rho}\big{)}$. As horizontal subspaces are invariant under right $G$-action, $[\delta_{(t)},e]_{\bar{\rho}}h(\delta_{(0)})$ is also a horizontal lift of $\delta_{(t)}$ on $P_{\bar{\rho}}$ passing through $[\delta_{(0)},e]_{\bar{\rho}}h(\delta_{(0)})=f\big{(}[\delta_{(0)},e]_{\rho}\big{)}$. Therefore, we have
$f\big{(}[\delta_{(t)},e]_{\rho}\big{)}=[\delta_{(t)},e]_{\bar{\rho}}h(\delta_{(0)})\,.$ (4.5)
Comparing (4.4)-(4.5), we have $h(\delta_{(t)})=h(\delta_{(0)})$ for all $t\in[0,1]$. This implies $h(\delta_{(t)})$ is equal to $h$ acting on the constant path at $a$. So $h$ is a constant, and therefore, $\rho$ and $\bar{\rho}$ are conjugate. ∎
@@
Let $\pi:P\to M$ be a principal $G$-bundle, and let $A$ be a cone-flat connection on $P$ with curvature $F_{A}=-\zeta B$. Then there exists $\xi\in\mathfrak{g}$ such that for any contractible loop $\gamma$ starting at $a\in M$ and any oriented disk $D$ in $M$ with $\partial D=\gamma$ ($\partial D$ and $\gamma$ also have the same orientation), the holonomy along $\gamma$ is
$hol(\gamma)=\mathrm{exp}\Big{[}\Big{(}\int_{D}\zeta\,\Big{)}\,\xi\Big{]}\,.$
@@
Take a point $u_{0}\in P$ on the fiber of $a\in M$. Consider the holonomy bundle
$\hat{P}=\{u\in P\,|\,\text{there exists a horizontal curve }\tilde{\delta}\text{ such that }\tilde{\delta}(0)=u_{0},\tilde{\delta}(1)=u\}.$
$\hat{P}$ is a principal bundle over $M$, and its structure group $\hat{G}$ is the holonomy group of $P$ at $u_{0}$. Let $\hat{A}$ be the restriction of $A$ on $\hat{P}$. By the holonomy theorem of Ambrose-Singer [AS], the Lie algebra of $\hat{G}$ is
$\hat{\mathfrak{g}}=\mathrm{span}\left\{F_{\hat{A}}\!\left(v_{1}^{H},v_{2}^{H}\right)\big{|}\;v_{1}^{H},v_{2}^{H}\text{ are horizontal vectors at }u\text{ for some }u\in\hat{P}\right\}.$
By assumption, $F_{A}=-\zeta B$, or more precisely $F_{A}=-(\pi^{*}\zeta)B$. Since $B$ is covariantly constant, it is equal to some $\xi\in\hat{\mathfrak{g}}$ at any point in $\hat{P}$. Hence, $F_{\hat{A}}\left(v_{1}^{H},v_{2}^{H}\right)\in\mathbb{R}\xi$. So $\hat{\mathfrak{g}}$ is 1-dimensional and abelian. Then $\hat{A}=-\theta\otimes\xi$ for some $\theta\in\Omega^{1}(\hat{P})$ and $d\theta=\hat{\pi}^{*}\zeta$, where $\hat{\pi}:\hat{P}\to M$ is the projection.
For an arbitrary contractible loop $\gamma$ and a disk $D$ such that $\partial D=\gamma$, there exists a contractible neighborhood $U\subset M$ of $D$. Then $\hat{P}|_{U}=U\times\hat{G}$ is trivial. Let $\sigma:U\to\hat{P}|_{U}$ be a local section and $\psi:\hat{P}|_{U}\to U\times\hat{G}$ be a trivialization such that $\psi\circ\sigma(p)=(p,e)$ for $p\in M$. Then $(\psi^{-1})^{*}\hat{A}=(-\sigma^{*}\theta\otimes\xi,0)+(0,MC_{\hat{G}})$, where $MC_{\hat{G}}:T\hat{G}\to\hat{\mathfrak{g}}$ is the Maurer-Cartan form of $\hat{G}$ sending a vector to the corresponding invariant vector field. Observe that $d(\sigma^{*}\theta)=\zeta$. The horizontal lift $\tilde{\gamma}$ of $\gamma$ with $\tilde{\gamma}(0)=\sigma(a)$ satisfies $\psi\circ\tilde{\gamma}(t)=(\gamma(t),g(t))$ with $g(t)\in\hat{G}$ and $g(0)=e$. With the horizontal vectors in the kernel space of the connection one-form, we have
@@
$=-(\sigma^{*}\theta)\big{(}\gamma^{\prime}(t)\big{)}\cdot\xi\Big{|}_{\gamma(t_{0})}+MC_{\hat{G}}\big{(}g^{\prime}(t)\big{)}\Big{|}_{g(t_{0})}.$
Hence, we find
@@
$hol(\gamma)=g(1)=\mathrm{exp}\Big{[}\Big{(}\int_{\gamma}\sigma^{*}\theta\,\Big{)}\,\xi\Big{]}=\mathrm{exp}\Big{[}\Big{(}\int_{D}\zeta\,\Big{)}\,\xi\Big{]}.$
Now we proceed to prove surjectivity.
@@
Let $\pi:P\to M$ be a principal $G$-bundle with a cone-flat connection $A$ with respect to $\zeta$. Then there exists a homomorphism $\rho:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G$ such that $(P_{\rho},A_{\rho})$ and $(P,A)$ are equivalent.